OpenStudy (anonymous):

How many cones of radius 1 m and height 1 m have total volume equal to the volume of a sphere with radius 1 m?

OpenStudy (raden):

solve for x x * 1/3 pi r^2 h = 4/3 pi r^3 with r=1 and h=1

OpenStudy (anonymous):

thanks

OpenStudy (raden):

you're welcome so, what's ur answer :)

OpenStudy (anonymous):

3

OpenStudy (raden):

incorrect, try again :)

OpenStudy (anonymous):

5

OpenStudy (raden):

no

OpenStudy (anonymous):

6?

OpenStudy (raden):

noproblem.. well, look at this how to get x x * 1/3 pi r^2 h = 4/3 pi r^3 subtitute the values of r=1 and h=1, will giving us x * 1/3 pi (1)^2 (1) = 4/3 pi (1)^3 x * 1/3 pi = 4/3 pi multiple by 3/pi from both sides, it can be x * 1/3 pi * 3/pi = 4/3 pi * 3/pi x = 4 so, u need 4 cones

OpenStudy (anonymous):

ohh thanks ha

OpenStudy (raden):

glad could help u ;)